A certain coil spring with an unstretch length of 1m requya force of 5N to stretch it 0.1cm. what work is done in stretching it by 1cm if the elastic limit is still not exceeded
1
Expert's answer
2019-01-21T11:14:10-0500

Calculate the work done by spring by the relation as follows:

E=(k〖(X)〗^2)/2

Here k is the spring constant and X is the stretching .

Calculate k by the relation as follows

k=F/x

Here F is the force applied and x is the displacement is spring.

k=(5 N)/(0.1 cm) = 5000 N/m

Substitute the value of k in energy relation.

E=((5000 N/m)〖(1 cm)〗^2)/2

E=((5000 N/m)〖(0.01 m)〗^2)/2

E=0.5/2= 0.25 J